Exam 3-Gases and Quantum Mechanics Question 23 Part A 113 21 mol of a gas occupies 56.2 L at 44°C and 793 torr, 4 62 mol of this gas occupies under these conditions L O 14.7 O 30.9 O 80.9 O 478 O 61.7 Submit Request Answer Provide Feedback
Answer #1

Given:

Vi = 56.2 L

ni = 3.21 mol

nf = 4.62 mol

use:

Vi/ni = Vf/nf

56.2 L / 3.21 mol = Vf / 4.62 mol

Vf = 80.9 L

Answer: 80.9 L